Choose the answer that correctly completes the statement: At 0∘C and 0.500 atm, 4.48 LNH3(g)

a) contains 0.20 mol NH3
b) has a mass of 3.40 g
c) contains 6.02×1022 molecules
d) contains 0.40 mol NH3
Answer #1

Given:

P = 0.5 atm

V = 4.48 L

T = 0.0 oC

= (0.0+273) K

= 273 K

find number of moles using:

P * V = n*R*T

0.5 atm * 4.48 L = n * 0.08206 atm.L/mol.K * 273 K

n = 0.1 mol

Molar mass of NH3,

MM = 1*MM(N) + 3*MM(H)

= 1*14.01 + 3*1.008

= 17.034 g/mol

use:

mass of NH3,

m = number of mol * molar mass

= 0.1 mol * 17.03 g/mol

= 1.70 g

Number of molecule = number of mol * Avogadro’s number

= 0.1 mol * 6.02*10^23 molecules/mol

= 6.02*10^22 molecule

Answer: c